Two cranes are used to lift a heavy concrete pillar. One crane pulls straight up with a force of 7200N. The other with a force of 5300N to the left at an angle of 34 degrees to the horizontal.
a) What is the magnitude of the combined force, to the nearest tenth of a newton?
b) At what angle does the pillar move to the nearest tenth of a degree?
Help please

Answer by htmentor(1343)

F1 = 7200 N straight up
F2 = 5300 N at 34 deg
To resultant force R = sqrt(Fx^2 + Fy^2), where Fx and Fy are the net forces
in the x and y directions.
Since one crane pulls straight up, the x force has one component:
Fx = F2*cos(34)
The y force has two components:
Fy = 7200 + F2*sin(34)
The angle the pillar moves is in the direction of the applied force:
A = arctan(Fy/Fx)